integrate the following integral
∫ sin^2 bx dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int sin^{2}bxdx
I = \frac{1}{2}\int 2sin^{2}bxdx
I = \frac{1}{2}\int (1-cos2bx)dx
I = \frac{1}{2}\int dx-\frac{1}{2}\int cos2bxdx
I = \frac{x}{2}-\frac{1}{2}\int cos2bxdx
I = \frac{x}{2}-\frac{1}{4b}sin2bx+c